### v3

@@ Abstract
Spatio temporal data consist of measurement for one or more raster fields such as weather, traffic volume, crime rate, or disease incidents. Advances in modern technology have increased the number of available information for this type of data hence the rise of multidimensional data. In this paper we take advantage of the multidimensional structure of the data but also its temporal and spatial structure. We will be using the NCAR Climate Data Gateway website which provides data discovery and access services for global and regional climate model data. The daily values of total precipitation (prec), maximum (tmax), and minimum (tmin) temperature are combined to create a multidimensional data called tensor (a multidimensional array). In this paper, we propose a spatio temporal principal component analysis to initialize CP decomposition component. We take full advantage of the spatial and temporal structure of the data in the initialization step for cp component analysis. The performance of our method is tested via comparison with most popular initialization method. We also run a clustering analysis to further show the performance of our analysis.
@@ 1 Introduction
Dimensionality-reduction methods offer an alternative pathway. Principal Component Analysis (PCA) [jolliffe2016principal, bishop2006pattern] also known in the climate sciences as Empirical Orthogonal Function analysis [HannachiEtal2007, MonahanEtal2009]—is the standard tool for deriving compact spatial and temporal representations. While PCA efficiently extracts dominant variance patterns, it treats variables independently and imposes orthogonality constraints that lack clear physical interpretation [HannachiEtal2007].
@@ 2.2 Spatial properties
Here our analysis focus on the spatial properties of the climate variables for each of the four continental seasons: winter (from December to January), spring (from March to May), summer (from June to August) and fall (from September to November). To emphasize spatial patterns, we look at maps of Spearman correlation coefficients. Figure 1 shows Spearman correlation coefficients computed with respect to the reference location indicated by the black square (same color scale across all panels). In all panels, correlations decrease with distance from the reference location. The correlation coefficients are mostly positive. The patterns for tmax and tmin are broadly similar, with slight differences in winter and summer compared to precipitation’s. tmax and tmin correlations remain strong in all seasons, with minor reductions in spatial extent during summer. Precipitation (left column) exhibits weaker and more spatially variable correlations compared to temperature.
@@ 3.1 Spatio-Temporal Principal Component Analysis
Principal component analysis (PCA) is a powerful and widely used method in multivariate statistics. Principal Component Analysis (PCA) [hotelling1933analysis] can be performed on the sample covariance matrix $\mathbf{S}=\tfrac{1}{i}\mathbf{X_{k}}^{\top}\mathbf{X_{k}}$, where $\mathbf{X_{k}}$ is an $I\times J$ centered data matrix (observations of $j$ variables on $i$ independent objects). However, this approach do not take into account the dependence among nearby locations, when applied to spatial (geographic) data. The eigenvalues are nonnegative, and the principal components are given by the eigenvectors associated with the positive eigenvalues. cliff1ord addressed this by introducing the concept of ‘spatial autocorrelation’. Incorporating this into PCA requires specifying a spatial-weights matrix $\mathbf{W}$ that encodes neighborhood or proximity relations among the spatial units under study. Several literature on spatial PCA (SPCA) discusses how to define $\mathbf{W}$ [Demšar01012013, Cartone03042021, 7846858]. For this paper, our focus will be on Spatio Temporal Principal Component Analysis (STPCA); see [Krzyśko02012024]. STPCA treats feature observations as functional time series, emphasizing their temporal evolution while preserving spatial structure. Its aim is to extract the dominant modes of variability in spatio-temporal data by jointly incorporating temporal dynamics, spatial relationships, and variation in feature magnitudes. To run an STPCA, a list of matrices is considered with the rows representing $i$ time steps and columns $j$ locations or region of values of the $k$ features/variables being considered. In fact, first the data matrix is transformed into a multivariate functional data set; see [gorecki2018selected, gorecki2012functional]. The transformation into a functional data is done by creating a matrix $A$ of size $I\times pk$ through fourier basis eq. 1 and keeping the coefficients.
$\begin{aligned} \phi_{0}(t)&=\frac{1}{\sqrt{T}},\\ \phi_{2b-1}(t)&=\sqrt{\frac{2}{T}}\;\sin\!\left(\frac{2\pi b\,t}{T}\right),\\ \phi_{2b}(t)&=\sqrt{\frac{2}{T}}\;\cos\!\left(\frac{2\pi b\,t}{T}\right),\end{aligned}\qquad\begin{aligned} &b=1,2,\ldots,\frac{B_{k}}{2},\\ &k=1,2,\ldots,m,\\ &t\in[0,n],\end{aligned}\quad\text{with }B_{k}\text{ even.}$ (1)
@@ 5 Conclusions
In this study, we employed a Spatio-Temporal Principal Component Analysis (STPCA) to initialize the factors in our CANDECOMP/PARAFAC (CP) decomposition, which was subsequently optimized using the Alternating Least Squares (ALS) algorithm. The numerical experiments yielded promising results, demonstrating a distinct advantage in leveraging the spatio-temporal dependencies inherent in the data into the CP factorization process. Model performance was assessed through the computation of the relative reconstruction error for each configuration. Furthermore, we applied k-means clustering to the extracted factor matrices and evaluated clustering quality using silhouette scores. The proposed STPCA+CPD approach consistently outperformed baseline methods, confirming the benefits of our initialization strategy. For future research, this framework could be extended by integrating deep learning architectures to enhance its capacity to capture complex spatio-temporal dynamics. Additionally, we plan to investigate more robust spatio-temporal tensor decomposition schemes and to generalize the tensor framework for forecasting and downscaling applications.
[…]
Supplementary information
[…]
AcknowledgementsThe author would like to thank Professor Julie Carreau for helpful discussion concerning this project.
Declarations
  • The author declares that no funds, grants, or other support were received during the preparation of this manuscript.
  • The author hereby declare that they have no known competing financial interests or personal relationships that could have potentially influenced the work presented in this paper.
  • All data is publicly available and cited.
  • No materials was used
  • Codes available are available upon request.
  • All work in this manuscript was done by the author.

### v4

@@ Abstract
Spatio temporal data consist of measurement for one or more raster fields such as weather, traffic volume, crime rate, or disease incidents. Advances in modern technology have increased the number of available information for this type of data hence the rise of multidimensional data. In this paper we take advantage of the multidimensional structure of the data but also its temporal and spatial structure. We will be using the NCAR Climate Data Gateway website which provides data discovery and access services for global and regional climate model data. The daily values of total precipitation (prec), maximum (tmax), and minimum (tmin) temperature are combined to create a multidimensional data called tensor (a multidimensional array). In this paper, we propose a spatio temporal principal component analysis to initialize CP decomposition component. We take full advantage of the spatial and temporal structure of the data in the initialization step for cp component analysis. The performance of our method is tested via comparison with most popular initialization methods. We also run a clustering analysis to further show the performance of our analysis.
@@ 1 Introduction
Dimensionality-reduction methods offer an alternative pathway. Principal Component Analysis (PCA) [Ye_2025, jolliffe2016principal, bishop2006pattern] also known in the climate sciences as Empirical Orthogonal Function analysis [HannachiEtal2007, MonahanEtal2009]—is the standard tool for deriving compact spatial and temporal representations. While PCA efficiently extracts dominant variance patterns, it treats variables independently and imposes orthogonality constraints that lack clear physical interpretation [HannachiEtal2007].
@@ 2.2 Spatial properties
Here our analysis focus on the spatial properties of the climate variables for each of the four continental seasons: winter (from December to January), spring (from March to May), summer (from June to August) and fall (from September to November). To emphasize spatial patterns, we look at maps of Spearman correlation coefficients. Figure 1 shows Spearman correlation coefficients computed with respect to the reference location indicated by the black square (same color scale across all panels). In all panels, correlations decrease with distance from the reference location. The patterns for tmax and tmin are broadly similar, with slight differences in winter and summer compared to precipitation’s. tmax and tmin correlations remain strong in all seasons, with minor reductions in spatial extent during summer. Precipitation (left column) exhibits weaker and more spatially variable correlations compared to temperature. The correlation are all positive.
@@ 3.1 Spatio-Temporal Principal Component Analysis
Principal component analysis (PCA) is a powerful and widely used method in multivariate statistics. Principal Component Analysis (PCA) [hotelling1933analysis] can be performed on the sample covariance matrix $\mathbf{S}=\tfrac{1}{i}\mathbf{X_{k}}^{\top}\mathbf{X_{k}}$, where $\mathbf{X_{k}}$ is an $I\times J$ centered data matrix (observations of $j$ locations on $i$ time steps). However, this approach do not take into account the dependence among nearby locations, when applied to spatial (geographic) data. The eigenvalues are nonnegative, and the principal components are given by the eigenvectors associated with the positive eigenvalues. cliff1ord addressed this by introducing the concept of ‘spatial autocorrelation’. Incorporating spatial autocorrelation chen2023spatial into PCA requires specifying a spatial-weights matrix $\mathbf{W}$ that encodes neighborhood or proximity relations among the spatial units under study. Several literature on spatial PCA (SPCA) discusses how to define $\mathbf{W}$ [Cartone03042021, 7846858]. For this paper, our focus will be on Spatio Temporal Principal Component Analysis (STPCA); see [Krzyko02012024]. STPCA treats feature observations as functional time series, emphasizing their temporal evolution while preserving spatial structure. Its aim is to extract the dominant modes of variability in spatio-temporal data by jointly incorporating temporal dynamics, spatial relationships, and variation in feature magnitudes. To run an STPCA, a list of matrices is considered with the rows representing $i$ time steps and columns $j$ locations or region of values of the $k$ features/variables being considered. In fact, first the data matrix is transformed into a multivariate functional data set; see [gorecki2018selected, gorecki2012functional]. The transformation into a functional data is done by creating a matrix $A$ of size $I\times pk$ through fourier basis eq. 1 and keeping the coefficients.
$\begin{aligned} \phi_{0}(t)&=\frac{1}{\sqrt{n}},\\ \phi_{2b-1}(t)&=\sqrt{\frac{2}{n}}\;\sin\!\left(\frac{2\pi b\,t}{n}\right),\\ \phi_{2b}(t)&=\sqrt{\frac{2}{n}}\;\cos\!\left(\frac{2\pi b\,t}{n}\right),\end{aligned}\qquad\begin{aligned} &b=1,2,\ldots,\frac{B_{k}}{2},\\ &k=1,2,\ldots,m,\\ &t\in[0,n],\end{aligned}\quad\text{with }B_{k}\text{ even.}$ (1)
@@ 5 Conclusions
In this study, we employed a Spatio-Temporal Principal Component Analysis (STPCA) to initialize the factors in our CANDECOMP/PARAFAC (CP) decomposition for better result using the Alternating Least Squares (ALS) algorithm. The numerical experiments yielded promising results, demonstrating a distinct advantage in leveraging the spatio-temporal dependencies inherent in the data into the CP factorization process. Model performance was assessed through the computation of the relative reconstruction error for each configuration. Furthermore, we applied k-means clustering to the extracted factor matrices and evaluated clustering quality using silhouette scores. The proposed STPCA+CPD approach consistently outperformed baseline methods, confirming the benefits of our initialization strategy. For future work, this framework could be extended by integrating deep learning architectures to enhance its capacity to capture complex spatio-temporal dynamics. Additionally, we plan to investigate more robust spatio-temporal tensor decomposition schemes and to generalize the tensor framework for forecasting and downscaling applications.
Acknowledgements
The author would like to thank Professor Julie Carreau for helpful discussion concerning this project.